OpenStudy (anonymous):

with an average acceleration of -1.2m/s2, how long will it take a cylist to bring a bicycle with an initial speed of 6.5/s to a complete stop?

random231 (random231):

acceleration (a)= -1.2 m/s^2 initial velocity(u)= 6.5 m/s using v=u+at 0=6.5-1.2t -6.5=-1.2t t=6.5/1.2 t=5.4111 seconds